OpenStudy (anonymous):

- 7xy^2 divide 28x^5y

OpenStudy (anonymous):

\[-7xy^2/28x^5y=-y/4x^4\] y^2/y = y^+1, so there's a y on top. x/x^5 = x^-4, so x^4 goes on the bottom. -7/28 simplifies to -1/4

OpenStudy (anonymous):

ty